OpenStudy (anonymous):

(3+2i)/(5+i)..? Totally stumped. Thanks everyone!

OpenStudy (saifoo.khan):

\[\frac{3+2i}{5+i} \times \frac{5-i}{5-i}\]

OpenStudy (anonymous):

Oh my god that's it! Totally forgot about conjugates, thank you so much!

OpenStudy (agreene):

\[\frac{17}{26}+\frac{7i}{26}\]